Examples Find the y- Parametes. 4 Ii 2 2ux iz MA + T + 12
Answer #1

then Va I in X Von 2v2 + H 12. 412 = Vz-Vat8u2 +VxV2 = 2Va uiz= Votta from a + 2 II = YuvI+Y12V2 Iz=% VitY22 V2 Applying nodai= But 3 + Vy 2 2 i = 1.581 +0.50 Y-parameters - 11.5 0.5 - 315 -15